# Efficient 3D Molecular Generation with Flow Matching and Scale Optimal Transport

### v2

@@ Abstract
Generative models for 3D drug design have gained prominence recently for their potential to design ligands directly within protein pockets. Current approaches, however, often suffer from very slow sampling times or generate molecules with poor chemical validity. Addressing these limitations, we propose Semla, a scalable E(3)-equivariant message passing architecture. We further introduce a molecular generation model, SemlaFlow, which is trained using flow matching along with scale optimal transport, a novel extension of equivariant optimal transport. Our model produces state-of-the-art results on benchmark datasets with just 100 sampling steps. Crucially, SemlaFlow samples high quality molecules with as few as 20 steps, corresponding to a two order-of-magnitude speed-up compared to state-of-the-art, without sacrificing performance. Furthermore, we highlight limitations of current evaluation methods for 3D generation and propose new benchmark metrics for unconditional molecular generators. Finally, using these new metrics, we compare our model’s ability to generate high quality samples against current approaches and further demonstrate SemlaFlow’s strong performance.
@@ 1 Introduction
Generative models for 3D drug design have recently seen a surge of interest due to their potential to design binders directly within protein pockets. Some recently proposed models have attempted to directly generate ligands within protein pockets [35, 15, 42]. More thorough analysis, however, revealed that many of these models generate ligands with unrealistic binding poses [17]. Others have attempted to train unconditional 3D molecular generators as a starting point [19, 47, 56, 20, 33, 58, 26]. Specifically, models which apply diffusion [18, 46] to molecular coordinates have been particular popular. However, these models also suffer significant practical limitations; namely, they almost all require hundreds or even thousands of forward passes during generation, making them impractical for most downstream applications. Many also generate chemically unrealistic or poor quality samples when applied to datasets of drug-like molecules.
@@
For molecular generators which represent molecules as strings or 2D graphs, fine-tuning for specific protein pockets has proven very fruitful [7, 31, 3] and is currently standard practice in the field. Frequently, these models are guided into optimised chemical spaces using reinforcement learning (RL). This approach, while very effective, requires that high quality molecules can be sampled very quickly. Existing 3D molecular generators, which use fully-connected message passing, exhibit very poor scaling to larger molecules and larger model sizes. State-of-the-art unconditional generators [56, 26] take minutes to sample a single batch, making them impractical for RL-based fine-tuning.
@@
In this work we tackle this problem from two directions. Firstly, we introduce a novel equivariant architecture for 3D molecular generation which exhibits significantly better efficiency and scalability than existing approaches. Even with 3 times as many parameters as the current state-of-the-art, our model processes a batch of molecules more than twice as quickly. Secondly, we introduce a novel type of optimal transport, scale optimal transport, which extends equivariant optimal transport [23, 47] to account for different sizes of molecules. Using our model architecture, we then train an unconditional molecular generator, SemlaFlow, using flow matching with scale optimal transport. With only 100 sampling steps our model outperforms all existing approaches on two benchmark datasets. We further show that with as few as 20 sampling steps SemlaFlow’s performance is comparable to the existing state-of-the-art while producing a 2-orders-of-magnitude speedup in sampling times. Finally, we highlight issues with current evaluation metrics and introduce two new benchmark metrics for unconditional molecular generation to combat these.
@@
We provide an overview of our molecular generation approach using scale optimal transport in Figure 1 and we summarise our key contributions as follows:
@@
  • Scale Optimal Transport – We achieve a significant reduction in the transport cost by extending equivariant optimal transport to account for molecular size.
  • SemlaFlow – A state-of-the-art molecular generator trained using flow matching, which provides up to a 70-fold improvement in sampling time compared to existing approaches.
  • Evaluation Metrics – We introduce new evaluation metrics for unconditional molecular generators which address a number of issues with current benchmarks.
Figure 1: An overview of our approach. Firstly, prior coordinate noise is sampled according to the respective scale of the molecules (atom and bond type noise is not shown for brevity). Our trained model, SemlaFlow, transports this noise to the data distribution to generate a molecular sample.
@@ 2 Background
Flow matching seeks to learn a generative process which transports samples from a noise distribution $p_{\mathrm{noise}}$ to samples from a data distribution $p_{\mathrm{data}}$. Conditional flow matching (CFM) has emerged in different flavors as an effective way to train flow matching models [2, 30, 28]. CFM works in a simulation-free manner by interpolating between noise and a data sample $x_{1}\sim p_{\mathrm{data}}(\mathbf{x}_{1})$. To do this a time-dependent conditional flow $p_{t|1}(\cdot|\mathbf{x}_{1})$ is defined, from which a conditional vector field $u_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})$ can be derived. Typically, a model $v_{t}^{\theta}(\mathbf{x}_{t})$ is trained to regress the vector field, but other formulations [10, 48] have trained a model to estimate the distribution $p_{1|t}^{\theta}(\cdot|\mathbf{x}_{t})$, which reconstructs clean data from noisy data. The vector field can then be constructed using the expectation:
@@
The static optimal transport (OT) problem seeks to identify a coupling $\pi$ between samples from two distributions that minimises a cost function [32]. Identifying the $\pi$ which is the 2-Wasserstein OT map defined by
$W^{2}_{2}=\inf_{\pi\in\Pi}\int c(x_{0},x_{1})\,\pi(\mathrm{d}x_{0},\mathrm{d}x_{1})$ (2)
where $c(x_{0},x_{1})=\lVert x_{0}-x_{1}\lVert_{2}$ is the cost function and $\Pi$ is the set of all couplings, has proven to be a successful strategy to improve the efficiency of CFM. Previous work has proposed mini-batched [36, 51] and group equivariant [23, 47] variations of this. In equivariant OT, the cost function is replaced by $\min_{g\in\mathcal{G}}\hat{c}(x_{0},x_{1})=\lVert x_{0}-T_{g}x_{1}\lVert_{2}$ where $T_{g}$ is a linear representation of $g\in\mathcal{G}$ [45].
@@
Köhler et al. [24] showed that, if a base density $p_{0}(\mathbf{x}_{0})$ is $\mathcal{G}$-invariant and a target density $p_{1}(\mathbf{x}_{1})$ is generated by following a $\mathcal{G}$-equivariant vector field, then $p_{1}(\mathbf{x}_{1})$ is also $\mathcal{G}$-invariant. We use this finding to ensure that the density of molecular coordinates learned by our model is $\mathcal{G}$-invariant by only applying equivariant updates and sampling coordinate noise from an isotropic Gaussian. For molecular generation we are concerned with the group $\mathcal{G}=\mathrm{E}(3)\times S_{N}$ where $\mathrm{E}(3)$ is the Euclidean group in 3 dimensions, encompassing translations, rotations and reflections, and $S_{N}$ is the symmetric group for a set with $N$ elements – the group of all possible permutations.
@@ 3 The Semla Architecture
Existing state-of-the-art models for 3D molecular generation [19, 56, 26] use fully-connected, multi-layer perceptron (MLP)-based message passing layers. However, the computational cost of such layers scales quadratically in both the feature dimension and the number of atoms. Consequently, these layers become a significant computational bottleneck when scaling to larger, drug-like molecules.
@@
To alleviate this problem we propose Semla - a scalable equivariant model which uses multi-head latent graph attention, where message passing is performed on compressed latent representations. This extension allows us to scale the dimensionality of the node features and the number of learnable model parameters without leading to prohibitory increases in computational cost. We illustrate the Semla architecture in Fig. 2 and further expand on each component below.
Figure 2: Architectural overview of a full Semla model (left) and one Semla layer (right).
@@
Similarly to previous approaches, Semla makes use of $\mathrm{E}(3)$ invariant and equivariant features. Enforcing group symmetry provides strong inductive biases and improves sample complexity [6, 50, 19]. However, unlike previous molecular generation models, Semla does not distinguish between molecular coordinates and equivariant feature vectors, but rather treats them as a single learnable representation, which we refer to as coordinate sets. Coordinate sets are analogous to the sets of 3D geometric equivariant vectors found in other equivariant architectures [21, 44, 25]. We argue that this representation has two key benefits over previous molecular generation approaches. Firstly, learnable coordinate sets provide much more expressivity than models which store only one set of coordinates per molecule [19, 56]. Additionally, a joint representation of equivariant features allows for a simpler update mechanism – we can simply apply linear projections (without bias) to create and update the coordinate sets while maintaining equivariance.
@@
To ensure stable training we use normalisation layers throughout the model. LayerNorm [5] is applied to invariant features, and, for equivariant features, we adapt the normalisation scheme from MiDi [56] to allow for multiple coordinate sets. We hypothesise that this allows coordinate sets to learn equivariant features of different length scales, which helps to circumvent the problem of molecules of different sizes being normalised to have the same average vector norm. We further extend the normalisation to ensure that coordinates are zero-centred. An additional zero-centering, which we apply at the end of the model, ensures that the learned density is translation invariant [13, 59]. For the remainder of this paper we will use $\phi_{inv}(\cdot)$ and $\phi_{equi}(\cdot)$ to refer to the normalisation functions for invariant and equivariant features, respectively.
@@
Throughout this paper we denote invariant and equivariant features for an atom $i$ as $\mathbf{h}_{i}\in\mathbb{R}^{D}$ and $\mathbf{x}_{i}\in\mathbb{R}^{S\times 3}$, respectively, where $D$ is the model dimension and $S$ is the number of coordinate sets. We also use $N$ to refer to the number of atoms in the molecule. To simplify the notation we assume that operations applied to $\mathbf{x}_{i}$ implicitly correspond to the concatenation of the results of the operation applied to individual vectors, unless we make the coordinate set explicit using a superscript. For example, the norm of $\mathbf{x}_{i}$ is implicitly applied as $\lVert\mathbf{x}_{i}\rVert=\left[\lVert\mathbf{x}_{i}^{1}\rVert,\lVert\mathbf{x}_{i}^{2}\rVert,\dots,\lVert\mathbf{x}_{i}^{S}\rVert\right]$.
@@ 3.1 Feature Feed-forward
$\mathbf{h}_{i}^{\text{ff}}=\mathbf{h}_{i}+\Phi_{\theta}(\mathbf{\tilde{h}}_{i},\lVert\phi_{equi}(\mathbf{x}_{i})\rVert)\hskip 25.0pt\mathbf{\tilde{h}}_{k}=\phi_{inv}(\mathbf{h}_{k})\hskip 25.0pt\mathbf{\tilde{x}}_{k}=\mathbf{W}_{\theta}^{1}\phi_{equi}(\mathbf{x}_{k})$ (3)
$\mathbf{x}_{i}^{\text{ff}}=\mathbf{x}_{i}+\mathbf{W}_{\theta}^{2}\left(\sum_{j=1}^{S}\mathbf{\tilde{x}}_{i}^{j}\otimes\Psi_{\theta}(\mathbf{\tilde{h}}_{i})\right)$ (4)
where $\Phi_{\theta}$ and $\Psi_{\theta}$ are learnable multi-layer perceptrons, $\mathbf{W}_{\theta}^{1}\in\mathbb{R}^{P\times S}$ and $\mathbf{W}_{\theta}^{2}\in\mathbb{R}^{S\times P}$ are learnable weight matrices with hyperparameter $P$, and $\otimes$ is the outer product. For simplicity we set $P=M$, the message size, which is defined below. Similarly to the transformer architecture [52], $\Phi_{\theta}$ linearly maps features to $4D$, applies a non-linearity (we use SiLU [12] throughout) and then maps back to the model dimension, $D$.
@@ 3.2 Equivariant Graph Attention
In this section, we introduce a novel attention mechanism for 3D graph structures. Like previously proposed attention mechanisms for equivariant architectures [40, 25, 27], our model computes pairwise messages using a multi-layer perceptron (MLP). These pairwise messages are then split in two and passed to attention mechanisms for invariant and equivariant features, respectively. We describe each of these components in more detail below.
@@
Pairwise messages are computed using a 2-layer MLP which combines invariant node features with pairwise dot products from the coordinate sets. Similarly to previous approaches we compute messages between all pairs of nodes in the graph. Unlike models such as EGNN [40], MiDi [56] and EQGAT [25], however, we first compress the invariant node features into a smaller latent space, with dimensionality $M$, using a learnable linear map. This reduces the computational complexity of the pairwise MLP from $\mathcal{O}(N^{2}D^{2})$ to $\mathcal{O}(N^{2}M^{2})$ where $M\ll D$, leading to a significant reduction in the compute and memory overhead of the MLP, especially on larger molecules. It also allows us to scale the size of the invariant node features, $D$, independently of $M$. Formally, messages between nodes $i$ and $j$, which are split into invariant and equivariant attention components, are computed as follows:
$(\mathbf{m}_{i,j}^{(inv)},\mathbf{m}_{i,j}^{(equi)})=\Omega_{\theta}(\mathbf{\tilde{h}}_{i},\mathbf{\tilde{h}}_{j},\mathbf{\tilde{x}}_{i}\cdot\mathbf{\tilde{x}}_{j})\hskip 25.0pt\mathbf{\tilde{h}}_{k}=\mathbf{W}_{\theta}^{3}\phi_{inv}(\mathbf{h}_{k}^{\text{ff}})\hskip 25.0pt\mathbf{\tilde{x}}_{k}=\phi_{equi}(\mathbf{x}_{k}^{\text{ff}})$ (5)
where $\Omega_{\theta}$ is the pairwise message MLP and $\mathbf{W}_{\theta}^{3}\in\mathbb{R}^{M\times D}$ is a learnable weight matrix.
@@
Once messages have been computed a softmax operation is applied to produce attention weights between pairs of nodes. These weights are then used to aggregate node features by taking a weighted average. Since the message vectors can, in general, be smaller than the node features, each scalar in the message vector attends to a fixed number of scalars within the node feature vectors. We note that this attention implementation generalises the attention mechanism found in EQGAT and related models such as the Point Transformer [61], where each scalar in the message attends to exactly one scalar in the node features, and is very closely related to the multi-head attention mechanism adopted in GAT [53, 9] and the Transformer [52]. We also make use of the recently proposed variance preserving aggregation mechanism [41], which corresponds to multiplying the attended vectors by weights $w_{i}^{k}$. Overall, our invariant feature attention is computed as:
$\alpha_{i,j}^{k}=\frac{\text{exp}(m_{i,j}^{k,(inv)})}{\sum_{j^{\prime}=1}^{N}\text{exp}(m_{i,j^{\prime}}^{k,(inv)})}\hskip 25.0pt\mathbf{\tilde{h}}_{i}=\mathbf{W}_{\theta}^{4}\phi_{inv}(\mathbf{h}_{i}^{\text{ff}})$ (6)
[…]
@@
Similarly to the invariant features, messages for the equivariant features are used to apply an attention-based update. The attention function applied here is line with previous work [40, 56, 25], however we extend it to allow for multiple coordinate sets. Notably, we also find the use of softmax normalisation on raw messages to be beneficial for overall model performance. Analogously to the invariant feature attention, we also apply variance preserving updates to the equivariant features. Semla attention for equivariant features is therefore defined as follows:
$\alpha_{i,j}^{k}=\frac{\text{exp}(m_{i,j}^{k,(equi)})}{\sum_{j^{\prime}=1}^{N}\text{exp}(m_{i,j^{\prime}}^{k,(equi)})}\hskip 25.0pt\mathbf{\hat{x}}_{i,j}=\frac{\mathbf{\tilde{x}}_{j}-\mathbf{\tilde{x}}_{i}}{\lVert\mathbf{\tilde{x}}_{j}-\mathbf{\tilde{x}}_{i}\rVert}\hskip 25.0pt\mathbf{\tilde{x}}_{k}=\mathbf{W}_{\theta}^{6}\phi_{equi}(\mathbf{x}_{k}^{\text{ff}})$ (8)
$\mathbf{x}_{k}^{\text{out}}=\mathbf{x}_{k}^{\text{ff}}+\mathbf{W}_{\theta}^{7}\Big{(}\left[w_{i}^{1}a_{i}^{1},\dots,w_{i}^{K}a_{i}^{K}\right]^{T}\Big{)}\hskip 25.0pta_{i}^{k}=\sum_{j=1}^{N}\alpha_{i,j}^{k}\hat{x}_{i,j}^{k}\hskip 25.0ptw_{i}^{k}=\sqrt{\sum_{j=1}^{N}\alpha_{i,j}^{k}}$ (9)

### v3

# SemlaFlow – Efficient 3D Molecular Generation with Latent Attention and Equivariant Flow Matching

@@ Abstract
Methods for jointly generating molecular graphs along with their 3D conformations have gained prominence recently due to their potential impact on structure-based drug design. Current approaches, however, often suffer from very slow sampling times or generate molecules with poor chemical validity. Addressing these limitations, we propose Semla, a scalable E(3)-equivariant message passing architecture. We further introduce an unconditional 3D molecular generation model, SemlaFlow, which is trained using equivariant flow matching to generate a joint distribution over atom types, coordinates, bond types and formal charges. Our model produces state-of-the-art results on benchmark datasets with as few as 20 sampling steps, corresponding to a two order-of-magnitude speedup compared to state-of-the-art. Furthermore, we highlight limitations of current evaluation methods for 3D generation and propose new benchmark metrics for unconditional molecular generators. Finally, using these new metrics, we compare our model’s ability to generate high quality samples against current approaches and further demonstrate SemlaFlow’s strong performance.
@@ 1 INTRODUCTION
Generative models for 3D drug design have recently seen a surge of interest due to their potential to design binders for given protein pockets. Some recently proposed models have attempted to directly generate ligands within rigid pockets (Peng et al., 2022; Guan et al., 2023; Schneuing et al., 2022). More thorough analysis, however, revealed that many of these models generate ligands with unrealistic binding poses (Buttenschoen et al., 2024; Harris et al., 2023). Others have attempted to train unconditional 3D molecular generators as a starting point (Hoogeboom et al., 2022; Song et al., 2023; Vignac et al., 2023; Hua et al., 2024; Morehead and Cheng, 2024; Xu et al., 2024; Le et al., 2024). Specifically, models which apply diffusion (Ho et al., 2020; Song et al., 2021) to molecular coordinates have been particular popular. However, these models also suffer significant practical limitations; namely, they almost all require hundreds or even thousands of forward passes during generation, making them impractical for most downstream applications. Many also generate chemically unrealistic or poor quality samples when applied to datasets of drug-like molecules.
Figure 1: An overview of the inference procedure for SemlaFlow. Noise is firstly sampled from the prior distribution over atom types, bond types and coordinates. We then integrate the vector field by applying our current prediction to the SemlaFlow model and taking a step in the direction of the model’s prediction. To improve readability noisy samples are displayed with bonds inferred based on the 3D coordinates.
@@
For molecular generators which represent molecules as strings or 2D graphs, fine-tuning for specific protein pockets has proven very fruitful (Blaschke et al., 2020; Loeffler et al., 2024; Atance et al., 2022) and is currently standard practice in the field. Frequently, these models are guided into optimised chemical spaces using reinforcement learning (RL). This approach, while very effective, requires that high quality molecules can be sampled very quickly. Existing 3D molecular generators, which use fully-connected message passing, exhibit poor scaling to larger molecules and larger model sizes – state-of-the-art unconditional generators (Vignac et al., 2023; Le et al., 2024) take minutes to sample a single batch, making them impractical for RL-based fine-tuning.
@@
In this work we tackle this problem from two directions. Firstly, we introduce a novel equivariant architecture for 3D graph generation, Semla, where attention between nodes is applied in a reduced latent space. Our architecture exhibits significantly better efficiency and scalability than existing approaches; even with twice as many parameters as the current state-of-the-art, our model processes a batch of molecules more than three times as quickly. Secondly, we propose a flow matching model for 3D molecular generation which learns a joint distribution over the molecular topology, atomic coordinates and formal charges. Our model, SemlaFlow, provides state-of-the-art results with as few as 20 sampling steps, corresponding to a 2-order-of-magnitude increase in sampling speed compared to existing models. Finally, we also highlight issues with frequently used evaluation metrics for unconditional 3D generation and propose the use of energy and strain energy for benchmarking unconditional 3D generative models.
@@
We provide an overview of our molecular generation approach in Figure 1 and we summarise our key contributions as follows:
@@
  • SemlaFlow – A state-of-the-art molecular generator trained using flow matching with equivariant optimal transport, which provides a more than 100-fold improvement in sampling time compared to existing approaches while maintaining state-of-the-art performance.
  • Benchmarking – We introduce new evaluation metrics to address a number of shortcomings with existing metrics for 3D molecular generation.
@@ 2 BACKGROUND
Flow matching seeks to learn a generative process which transports samples from a noise distribution $p_{\mathrm{noise}}$ to samples from a data distribution $p_{\mathrm{data}}$. Conditional flow matching (CFM) has emerged in different flavors as an effective way to train flow matching models (Albergo and Vanden-Eijnden, 2023; Liu et al., 2023; Lipman et al., 2023). CFM works in a simulation-free manner by interpolating between noise and a data sample $x_{1}\sim p_{\mathrm{data}}(\mathbf{x}_{1})$. To do this a time-dependent conditional flow $p_{t|1}(\cdot|\mathbf{x}_{1})$ is defined, from which a conditional vector field $u_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})$ can be derived. Typically, a model $v_{t}^{\theta}(\mathbf{x}_{t})$ is trained to regress the vector field, but other formulations (Campbell et al., 2024; Stark et al., 2024a) have trained a model to estimate the distribution $p_{1|t}^{\theta}(\cdot|\mathbf{x}_{t})$, which reconstructs clean data from noisy data. The vector field can then be constructed using the expectation:
@@
Köhler et al. (2020) showed that, if a base density $p_{0}(\mathbf{x}_{0})$ is $\mathcal{G}$-invariant and a target density $p_{1}(\mathbf{x}_{1})$ is generated by following a $\mathcal{G}$-equivariant vector field, then $p_{1}(\mathbf{x}_{1})$ is also $\mathcal{G}$-invariant. We use this finding to ensure that the density of molecular coordinates learned by our model is $\mathcal{G}$-invariant by only applying equivariant updates and sampling coordinate noise from an isotropic Gaussian. For molecular generation we are concerned with the group $\mathcal{G}=\mathrm{E}(3)\times S_{N}$ where $\mathrm{E}(3)$ is the Euclidean group in 3 dimensions, encompassing translations, rotations and reflections, and $S_{N}$ is the symmetric group for a set with $N$ elements – the group of all possible permutations.
@@ 3 The Semla Architecture
Existing state-of-the-art models for 3D molecular generation (Hoogeboom et al., 2022; Vignac et al., 2023; Le et al., 2024) use fully-connected, multi-layer perceptron (MLP)-based message passing layers. However, the computational cost of such layers scales quadratically in both the feature dimension and the number of atoms. Consequently, these layers become a significant computational bottleneck when scaling to larger, drug-like molecules.
@@
To alleviate this problem we propose Semla - a scalable equivariant model which uses multi-head latent graph attention, where message passing is performed on compressed latent representations. This extension allows us to scale the dimensionality of the node features and the number of learnable model parameters without leading to prohibitory increases in computational cost. We illustrate one Semla architecture in Figure 2 and further expand on each component below.
Figure 2: Architectural overview of one Semla layer.
@@
Similarly to previous approaches, Semla makes use of both $\mathrm{E}(3)$ invariant and equivariant features. Enforcing group symmetry provides strong inductive biases and improves sample complexity (Bietti et al., 2021; Tahmasebi and Jegelka, 2024; Hoogeboom et al., 2022). However, unlike many existing molecular generation architectures, Semla does not distinguish between molecular coordinates and equivariant feature vectors, but rather treats them as a single learnable representation. We argue that this representation has two key benefits over previous molecular generation approaches. Firstly, sets of learnable equivariant features provide much more expressivity than models which store only one set of coordinates per molecule (Satorras et al., 2021; Hoogeboom et al., 2022; Vignac et al., 2023). Additionally, a joint representation of equivariant features allows for a simpler update mechanism than that proposed in EQGAT (Le et al., 2022) – we can simply apply linear projections (without bias) to create and update the feature vectors while maintaining equivariance.
@@
To ensure stable training we use normalisation layers throughout the model. LayerNorm (Ba et al., 2016) is applied to invariant features, and, for equivariant features, we adapt the normalisation scheme from MiDi (Vignac et al., 2023) to allow for multiple equivariant feature vectors. We hypothesise that this allows the model to learn equivariant features of different length scales, which helps to circumvent the problem of molecules of different sizes being normalised to have the same average vector norm. We further extend the normalisation to ensure that coordinates are zero-centred. An additional zero-centering, which we apply at the end of the model, ensures that the learned density is translation invariant (Garcia Satorras et al., 2021; Xu et al., 2022). For the remainder of this paper we will use $\phi_{inv}(\cdot)$ and $\phi_{equi}(\cdot)$ to refer to the normalisation functions for invariant and equivariant features, respectively. Both LayerNorm and our coordinate normalisation module contain a small number of learnable parameters, these are not shared between layers.
@@
We will denote invariant and equivariant features for an atom $i$ as $\mathbf{h}_{i}\in\mathbb{R}^{d_{inv}}$ and $\mathbf{x}_{i}\in\mathbb{R}^{d_{equi}\times 3}$, respectively, where $d_{inv}$ is the number of invariant scalar features and $d_{equi}$ is the number of equivariant feature vectors. We also use $N$ to refer to the number of atoms in the molecule. To simplify the notation, we assume that operations applied to $\mathbf{x}_{i}$ implicitly correspond to the concatenation of the results of the operation applied to individual vectors, unless we make the equivariant feature vector explicit using a superscript. For example, the norm of $\mathbf{x}_{i}$ is implicitly applied as $\lVert\mathbf{x}_{i}\rVert=\left[\lVert\mathbf{x}_{i}^{1}\rVert,\lVert\mathbf{x}_{i}^{2}\rVert,\dots,\lVert\mathbf{x}_{i}^{d_{equi}}\rVert\right]$.
@@ 3.1 Feature Feed-forward
$\mathbf{\tilde{h}}_{k}=\phi_{inv}(\mathbf{h}_{k})\hskip 25.0pt\mathbf{\tilde{x}}_{k}=\mathbf{W}_{\theta}^{1}\phi_{equi}(\mathbf{x}_{k})$ (2)
$\mathbf{h}_{i}^{\text{ff}}=\mathbf{h}_{i}+\Phi_{\theta}(\mathbf{\tilde{h}}_{i},\lVert\phi_{equi}(\mathbf{x}_{i})\rVert)$ (3)
$\mathbf{x}_{i}^{\text{ff}}=\mathbf{x}_{i}+\mathbf{W}_{\theta}^{2}\left(\sum_{j=1}^{d_{equi}}\mathbf{\tilde{x}}_{i}^{j}\otimes\Psi_{\theta}(\mathbf{\tilde{h}}_{i})\right)$ (4)
where $\Phi_{\theta}$ and $\Psi_{\theta}$ are learnable multi-layer perceptrons, $\mathbf{W}_{\theta}^{1}\in\mathbb{R}^{d_{equi}\times d_{equi}}$ and $\mathbf{W}_{\theta}^{2}\in\mathbb{R}^{d_{equi}\times d_{equi}}$ are learnable weight matrices, and $\otimes$ is the outer product. Similarly to the transformer architecture (Vaswani et al., 2017), $\Phi_{\theta}$ linearly maps features to $4\times d_{inv}$, applies a non-linearity (we use SiLU (Elfwing et al., 2018) throughout) and then maps back to $d_{inv}$.
@@ 3.2 Equivariant Graph Attention
In this section, we introduce a novel attention mechanism for 3D graph structures. Like previously proposed attention mechanisms for equivariant architectures (Satorras et al., 2021; Le et al., 2022; Liao and Smidt, 2023), our model computes pairwise messages using a multi-layer perceptron (MLP). These pairwise messages are then split in two and passed to attention mechanisms for invariant and equivariant features, respectively. We describe each of these components in more detail below.
@@
Pairwise messages are computed using a 2-layer MLP which combines invariant node features with pairwise dot products from the equivariant features. Similarly to previous approaches we compute messages between all pairs of nodes in the graph. Unlike models such as EGNN (Satorras et al., 2021), MiDi (Vignac et al., 2023) and EQGAT (Le et al., 2022), however, we first compress the invariant node features into a smaller latent space, with dimensionality $d_{l}$, using a learnable linear map. This reduces the computational complexity of the pairwise MLP from $\mathcal{O}(N^{2}d_{inv}^{2})$ to $\mathcal{O}(N^{2}d_{l}^{2})$ where $d_{l}\ll d_{inv}$, leading to a significant reduction in the compute and memory overhead of the MLP, especially on larger molecules. It also allows us to scale the size of the invariant node features independently of the node latent dimension. In Appendix A we provide experiments varying the size of this parameter and show that it is possible to produce a significant reduction in inference time with negligible drop in generation quality. Formally, messages between nodes $i$ and $j$, which are split into invariant and equivariant attention components, are computed as follows:
$(\mathbf{m}_{i,j}^{(inv)},\mathbf{m}_{i,j}^{(equi)})=\Omega_{\theta}(\mathbf{\tilde{h}}_{i},\mathbf{\tilde{h}}_{j},\mathbf{\tilde{x}}_{i}\cdot\mathbf{\tilde{x}}_{j})$ (5)
$\mathbf{\tilde{h}}_{k}=\mathbf{W}_{\theta}^{3}\phi_{inv}(\mathbf{h}_{k}^{\text{ff}})\hskip 25.0pt\mathbf{\tilde{x}}_{k}=\phi_{equi}(\mathbf{x}_{k}^{\text{ff}})$ (6)
where $\Omega_{\theta}$ is the pairwise message MLP and $\mathbf{W}_{\theta}^{3}\in\mathbb{R}^{d_{l}\times d_{inv}}$ is a learnable latent projection matrix.
@@
Once messages have been computed a softmax operation is applied to produce attention weights between pairs of nodes. These weights are then used to aggregate node features by taking a weighted average. Since the message vectors can, in general, be smaller than the node features, each scalar in the message vector attends to a fixed number of scalars within the node feature vectors. We note that this attention implementation generalises the attention mechanism found in EQGAT and related models such as the Point Transformer (Zhao et al., 2021), where each scalar in the message attends to exactly one scalar in the node features, and is very closely related to the multi-head attention mechanism adopted in GAT (Veličković et al., 2018; Brody et al., 2021) and the Transformer (Vaswani et al., 2017). We also make use of the recently proposed variance preserving aggregation mechanism (Schneckenreiter et al., 2024), which corresponds to multiplying the attended vectors by weights $w_{i}^{k}$. Overall, our invariant feature attention is computed as:
$\alpha_{i,j}^{k}=\frac{\text{exp}(m_{i,j}^{k,(inv)})}{\sum_{j^{\prime}=1}^{N}\text{exp}(m_{i,j^{\prime}}^{k,(inv)})}\hskip 15.0ptw_{i}^{k}=\sqrt{\sum_{j=1}^{N}(\alpha_{i,j}^{k})^{2}}$ (7)
$\mathbf{\tilde{h}}_{i}=\mathbf{W}_{\theta}^{4}\phi_{inv}(\mathbf{h}_{i}^{\text{ff}})\hskip 25.0pt\mathbf{a}_{i}^{k}=\sum_{j=1}^{N}\alpha_{i,j}^{k}\mathbf{\tilde{h}}_{j}^{k}$ (8)
$\mathbf{h}_{i}^{\text{out}}=\mathbf{h}_{i}^{\text{ff}}+\mathbf{W}_{\theta}^{5}\Big{(}\Big{\|}_{k=1}^{K}w_{i}^{k}\mathbf{a}_{i}^{k}\Big{)}$ (9)
@@
Similarly to the invariant features, messages for the equivariant features are used to apply an attention-based update. The attention function applied here is line with previous work (Satorras et al., 2021; Vignac et al., 2023; Le et al., 2022), however we extend it to allow for multiple equivariant feature vectors. Notably, we also find the use of softmax normalisation on raw messages to be beneficial for overall model performance. Analogously to the invariant feature attention, we also apply variance preserving updates to the equivariant features. Semla attention for equivariant features is therefore defined as follows:
$\mathbf{\hat{x}}_{i,j}=\frac{\mathbf{\tilde{x}}_{j}-\mathbf{\tilde{x}}_{i}}{\lVert\mathbf{\tilde{x}}_{j}-\mathbf{\tilde{x}}_{i}\rVert}\hskip 25.0pt\mathbf{\tilde{x}}_{k}=\mathbf{W}_{\theta}^{6}\phi_{equi}(\mathbf{x}_{k}^{\text{ff}})$ (10)
$\alpha_{i,j}^{k}=\frac{\text{exp}(m_{i,j}^{k,(equi)})}{\sum_{j^{\prime}=1}^{N}\text{exp}(m_{i,j^{\prime}}^{k,(equi)})}$ (11)
[…]
$\mathbf{x}_{k}^{\text{out}}=\mathbf{x}_{k}^{\text{ff}}+\mathbf{W}_{\theta}^{7}\Big{(}\left[w_{i}^{1}a_{i}^{1},\dots,w_{i}^{K}a_{i}^{K}\right]^{T}\Big{)}$ (13)